OpenStudy (anonymous):

What is the distance between the points (–8, –2) and (–7, –2)? A. 1 B. 6 C. 10 D. 15

OpenStudy (usukidoll):

we can just use the distance formula...

OpenStudy (usukidoll):

for .. \[(x_1,y_1) = (-8,-2) ..... (x_2,y_2) = (-7,-2)\]

OpenStudy (usukidoll):

distance formula is given as \[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (usukidoll):

we can just plug the points in for \[x_1 = -8, x_2 = -7, y_1 = -2, y_2 = -2\]

OpenStudy (usukidoll):

\[d=\sqrt{(-7-(-8))^2+(-2-(-2))^2}\] \[d=\sqrt{(-7+8)^2+(-2+2)^2}\] \[d=\sqrt{1^2+0^2}\] \[d=\sqrt{1+0} = \sqrt{1} = 1\]

OpenStudy (usukidoll):

the distance is 1

jimthompson5910 (jim_thompson5910):

Since the y coordinates are the same (both -2) you can also subtract the x coordinates and take the absolute value of that difference |x1-x2| = |-8 - (-7)| = |-8 + 7| = |-1| = 1 and you'll get the same answer. It's probably better to practice with the distance formula since that can be used for any two points.

OpenStudy (usukidoll):

distance formula comes back in vector calculus... I would use that.